150 likes | 305 Views
TAU - status. August 17. Status. Large panel (VP1) is working and producing fast clean signals. With 1, 2 or 4 HV channels signals look and behave exactly the same. First measurements of response to radiation – both electrons and muons.
E N D
TAU - status August 17
Status • Large panel (VP1) is working and producing fast clean signals. • With 1, 2 or 4 HV channels signals look and behave exactly the same. • First measurements of response to radiation – both electrons and muons. • Small panel (BN01) was tested and found to hold vacuum, should be tested for signals tomorrow.
Our hodoscope • Roughly 2.5 cm2 of active area. • Overnight measurement of 13hours beneath the panel give 1509 cosmic triggers • This corresponds to ~2muons per minute
Signal shape study (VP1) set-up • Quench resistors: 20 Mohm (per line) • 4 RO lines: voltage divider (100:1 Ohm) • MDT gas (Ar-CO2 93%-7%) 600 Torr • RO lines connected to scope 1 Ohm Ni – RO line 100 Ohm For each RO line
Efficiency tests Set-up Source Panel Hodoscope AND between the scintillators OR between the RO lines AND counter counter counter • Quench resistors: 20 Mohm(per line) • 4 RO lines: voltage divider (100:1 Ohm) • Every RO line into discriminator -> logic
First attempt – Voltage scan, MDT gas @600Torr • More triggers as Voltage rises!!! • -> BAD grounding
Adjustments • Every ground in the system was connected to the other grounds with a thick wire. • Coincidents between panel & trigger almost stopped • We raised the voltage without improvement • White discoloration of the active pixels is visible. • Gas was changed to the same gas type and pressure. • New working area in the panel. • Hodoscope was raised closer to the panel
Second attempt – Voltage scan, MDT gas @600Torr • Trigger is constant regardless to panel Voltage • Almost no coincident between trigger and signal
Background • Without the source sitting on the panel there is almost no background. • No coincident between signal and the trigger At all
Efficiency tests Set-up 2 Source Panel Hodoscope AND between the scintillators OR between the RO lines AND counter counter counter • Quench resistors: 20 Mohm(per line) • 4 RO lines: voltage divider (100:1 Ohm) • Every RO line into discriminator -> logic
Third attempt – Voltage scan, MDT gas @600Torr • Rate of upper scintillation counter is ~15KHz • Rate of trigger is ~25 • We believe that upper scintillation pad is at it’s rate limit
Plans for next week • Signal shape studies with BN01 panel • Rates study - continue with PV1, start with BN01 panel • Expand the study to CF4